OpenStudy (anonymous):

how can I determine if the following function is injective, sobrejective, or biyective. f(x)=(x-1)/(x-2) Can you show the work please

myininaya (myininaya):

injective is easy to tell does it pass the horizontal line test?

OpenStudy (anonymous):

I can figure out how the graph is mentally. however, I need to do it mathematically

myininaya (myininaya):

oh i see

OpenStudy (anonymous):

for example in one exam, how can I prove that that function is one of the above menciones

myininaya (myininaya):

so you want to just do it purely my definition

OpenStudy (anonymous):

yeah. havent found exercises like that

myininaya (myininaya):

we know it is 1 to 1 so we cannot provide a counterexample so proving it by definition...

OpenStudy (anonymous):

let see what raheen has to tell us

myininaya (myininaya):

we know it isn't surjective

myininaya (myininaya):

so its not bijective

OpenStudy (anonymous):

for the injective test just plug an arbitrary x (from the domain) for example x= 3 ===> y = 2/1 = 2 . now let y =2 and solve for x ( if you get the same x i.e x=3, ===> 2=(x-1)(x-2) ==> x= 3 ,then it's injective

myininaya (myininaya):

counterexample for surjective find inverse so x=(2y-1)/(y-1) y cannot be 1

myininaya (myininaya):

or i mean y is never 1

myininaya (myininaya):

not every real number was hit

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

from the injective test and the surjective test we know if the function is bijective or not.

myininaya (myininaya):

injective: Let a, b be real numbers such that f(a)=f(b) => (a-1)/(a-2)=(b-1)/(b-2) this is only true when a=b so this implies f(x)=(x-1)/(x-2) is injective

OpenStudy (anonymous):

is is special case then?

myininaya (myininaya):

example of a non 1-1 function is f(x)=x^2 since: let a,b be real numbers such that f(a)=f(b)=> a^2=b^2 this is true when a=b or when a=-b so since we have a being two different numbers then f(x)=x^2 is not 1 to 1